OpenStudy (anonymous):

x-y is a rational number Is this reflexive, symmetric, transitive, or anti symmetric

OpenStudy (mertsj):

None of the above

OpenStudy (anonymous):

Sorry but my professor has a different answer than you.. im just trying to figure out how it's transitive and how its symmetric

OpenStudy (mertsj):

You need to give more information then.

OpenStudy (mertsj):

Are x and y rational numbers?

OpenStudy (anonymous):

oh sorry they are all real numbers

OpenStudy (mertsj):

So could you please post the entire problem exactly as it is written?

OpenStudy (anonymous):

Determine whether the relation R on the set of all real numbers is reflexive, anti-symmetric and/or transitive, where x -y is a rational number

OpenStudy (mertsj):

And what is the relation R?

OpenStudy (anonymous):

the set of all real numbers

OpenStudy (mertsj):

I don't think that can be right because a relation is a set of ordered pairs. So it can be the set of real numbers.

OpenStudy (anonymous):

Yeah and y and x are in that relation and the values are elements of the set of all real numbers

OpenStudy (mertsj):

So, I guess the relation is x-y. So a rational number is equal to itself, so yes, it is reflexive. If x-y=b then b =x-y so it is symmetric. If x-y=a and b = c then x-y=c so it is also transitive.

OpenStudy (anonymous):

it couldnt be reflexive if x - y = to rational number because for it to be reflexive (a,a) and 1 -1 is not a = a.. as for being symmetric (a,b) is in the element R and (b,a) is in the element R however i dont get how its transitive

OpenStudy (mertsj):

I guess I don't understand the interpretation then. If x-y is a rational number, a rational number is certainly equal to itself.

OpenStudy (anonymous):

yes but if the rational number is x - y if x is 1 and y is 1... then how is negative one equal to 1 ?

OpenStudy (mertsj):

It isn't but if x is 1 and y is 1 then x-y = 0 and 0 is certainly equal to 0. But as I said, I must be mis-intrepreting the problem.

Directrix (directrix):

@Sceptre12 Perhaps the information here will help. http://tinyurl.com/b6sxsl9

OpenStudy (anonymous):

Thank you this definition helps alot